four times the certain number, minus 5, is equal to 7 more than twice the number. What is the number?
Answer by apachemexican1(64)

The equation is written out like this
4n-5=7+2n
so put all like terms together
4n-2n=7+5
then add or subtract
2n=12
then divide both sides by 2
2n/2=12/2
n=6
good luck!